# Chaotic Noncoherent SWIPT in Multi-Functional RIS-Aided Systems

### v1

@@ IV-C Transmit-Reflect-Harvest Trade-off Characterization
=$=\left\{\left({\rm SR}_{t},{\rm SR}_{r},P_{\rm EH}\right):{\rm SR}_{X}\geq{\rm SR}_{X,0},P_{\rm EH}\geq E_{\rm req}\right\},$ (18)
@@ V Numerical Results
We consider a transmission power $P_{\rm t}=30$ dBm, path loss at one meter distance $C_{0}=10^{-3.53}$, and the noise power at the MF-RIS, $U_{t}$, and $U_{r}$ is $-90$ dBm. We have the path loss exponent $\alpha_{sr}=\alpha_{rd_{t}}=\alpha_{rd_{r}}=3$ and a two-tap Rayleigh fading wireless channel with $\Omega_{\alpha,1}=\Omega_{\beta,t,1}=\Omega_{\beta,r,1}=0.8$, and $\Omega_{\alpha,2}=\Omega_{\beta,t,2}=\Omega_{\beta,r,2}=0.2$. The parameters for the considered non-linear EH model are: $\eta_{1}=0.9207\times 10^{3},\eta_{2}=0.0052\times 10^{9},$ and $R_{L}=5000$ $\Omega$ [11].
@@
Fig. 1 shows the impact of the reference length and the amplification factors on the SR performance with $\left(N_{t}=10,d_{rd_{t}}=25\>{\rm m}\right)$ and $\left(N_{r}=20,d_{rd_{r}}=30\>{\rm m}\right)$, respectively. By considering $\beta=60$, we vary $\phi$ to investigate its impact on ${\rm SR}_{X}$ for various $\Upsilon_{X}$. We observe that irrespective of $\Upsilon_{X}$, ${\rm SR}_{X}$ monotonically increases with $\phi$; the rate of increase in ${\rm SR}_{X}$ is initially linear but it saturates gradually. Also, we note a significant improvement in the ${\rm SR}_{X}$ performance with higher $\Upsilon_{X}$; for example, observe the performance gap at $\phi=30$ between ${\rm SR}_{r}$ for $\Upsilon_{r}=1$ and $\Upsilon_{r}=3$. However, this enhancement in ${\rm SR}$ performance comes at the cost of a higher $E_{\rm req}$. In other words, (16) shows that a higher $\Upsilon_{X}$ results in a higher $E_{\rm req}$, i.e., more energy is required at the MF-RIS to attain self-sustainability. Lastly, with $\Upsilon_{t}=\Upsilon_{r}=1$, the BER performance of MF-RIS merges with that of the element splitting STAR-RIS [9].
Figure 1: Impact of $\phi,\Upsilon_{t}$ and $\Upsilon_{r}$ on the ${\rm SR}$.
@@
Fig. 2 illustrates the proposed characterization with $N=120,N_{t}=20,N_{r}=40,N_{h}=60,d_{sr}=9$ m, $d_{rd_{t}}=20$ m, $d_{rd_{r}}=24$ m, and spreading factor $\beta=60$. Specifically, for the above stated set of system parameters, we vary the reference length $\phi$ such that $\zeta=\frac{\beta}{\phi}\in\mathbb{Z}$ and investigate its combined impact on ${\rm SR}_{t},{\rm SR}_{r},$ and $P_{\rm EH}$, respectively. We observe that, for a given $(N_{t},N_{r},N_{h})$ combination, having $\Upsilon_{t}>1$ and $\Upsilon_{r}>1$ results in significant enhancement of the proposed generalized ${\rm SR}-P_{\rm EH}$ region. Moreover, the figure demonstrates the inevitable role of transmit waveform design on the system performance. Conventional DCSK, i.e., $\phi=\beta$ results in the best IT and IR performance but a very poor EH performance, which implies a very high $N_{h}^{\min}$ in (IV-C). On the contrary, $\phi=1$ results in a completely opposite performance, i.e., worst IT and IR performance but excellent EH performance. Hence, depending on the application specific acceptable BER, we decide on the value of $\phi$ for a given $\beta$.
VI Conclusion
@@
We have proposed a DCSK-based MF-RIS-aided noncoherent SWIPT framework. Specifically, the MF-RIS elements are partitioned into three sub-surfaces, namely, IT, IR, and EH. While the IT and IR section serve the users in the transmission and reflection region, respectively, the EH section harvests energy from the incident signal to meet their energy requirement. Accordingly, we derived the BER in closed-form and characterized the harvested energy-BER trade-off. We also investigated the impact of this trade-off, i.e., appropriate choice of the amplification factors and the reference length $\phi$, on the problem of transmit waveform design.
@@
In order to evaluate ${\rm BER}_{X}$, we need to obtain the mean and variance of $\lambda_{p}^{(X)}$ from (III-A).
@@
$\mathbb{E}\{\lambda_{p}^{(X)}\}$ $=\frac{\beta\gamma_{X,0}}{\left(\beta+\phi\right)}N_{0}d_{p}\Lambda_{X,1}$
${\rm var}\{\lambda_{p}^{(X)}\}$ $=\frac{\beta N_{0}^{2}}{2}\left(\Lambda_{X,1}\frac{\gamma_{X,0}\left(\zeta+1\right)}{\beta+\phi}+\frac{1}{2}\right)$
$\!\!\!\!\!\!\!\!+\frac{\beta\Psi_{X}^{2}N_{0}^{2}}{4}\Lambda_{X,2}\left(\left(\zeta+1\right)+\frac{4\zeta\gamma_{X,0}}{\left(\beta+\phi\right)}\Lambda_{X,1}\right).$ (19)
@@
By using the Gaussian approximation [6] and assuming equally probable transmission of $d_{p}=\pm 1$, the system IT/IR BER based on the channel conditions is obtained as
@@
By replacing (VI) in (VI) and followed by generalization over the considered channel model, we obtain
[…]
Figure 2: Impact of system parameters on the ${\rm SR}-P_{\rm EH}$ region.

### v2

@@ IV Waveform Design Investigating Transmit-Reflect-Harvest Trade-off
$=\left\{\left({\rm SR}_{t},{\rm SR}_{r},P_{\rm EH}\right):{\rm SR}_{X}\geq{\rm SR}_{X,0},P_{\rm EH}\geq E_{\rm req}\right\},$ (15)
@@ V Numerical Results and Conclusion
We consider a transmission power $P_{\rm t}=30$ dBm, path loss at one meter distance $C_{0}=10^{-3.53}$, and the noise power at the MF-RIS, $U_{t}$, and $U_{r}$ is 90-90 dBm. We have the path loss exponent $\alpha_{sr}=\alpha_{rd_{t}}=\alpha_{rd_{r}}=3$ and a two-tap Rayleigh fading wireless channel with $\Omega_{\alpha,1}=\Omega_{\beta,t,1}=\Omega_{\beta,r,1}=0.8$, and $\Omega_{\alpha,2}=\Omega_{\beta,t,2}=\Omega_{\beta,r,2}=0.2$. The parameters for the considered non-linear EH model are: $\eta_{1}=0.9207\times 10^{3},\eta_{2}=0.0052\times 10^{9},$ and $R_{L}=5000$ $\Omega$ [14].
@@
Fig. 1(a) shows the impact of $\phi$ and the amplification factors $\Upsilon_{X}$ on the SR performance. By considering $\beta=60$ and the system parameters as stated in the figure, we vary $\phi$ to investigate its impact on ${\rm SR}_{X}$ for various $\Upsilon_{X}$. We observe that irrespective of $\Upsilon_{X}$, ${\rm SR}_{X}$ monotonically increases with $\phi$; the rate of increase in ${\rm SR}_{X}$ is initially linear but it saturates gradually. Also, we note a significant improvement in the ${\rm SR}_{X}$ performance with higher $\Upsilon_{X}$; for example, observe the performance gap at $\phi=30$ between ${\rm SR}_{r}$ for $\Upsilon_{r}=1$ and $\Upsilon_{r}=3$. However, this enhancement in $\rm SR$} performance comes at the cost of a higher $E_{\rm req}$. In other words, (13) shows that a higher $\Upsilon_{X}$ results in a higher $E_{\rm req}$, i.e., more energy is required at the MF-RIS to attain self-sustainability. Lastly, with $\Upsilon_{t}=\Upsilon_{r}=1$, the BER performance merges with that of the element splitting STAR-RIS [11]. Fig. 1(b) illustrates that for identical system parameters, SR decreases with increasing MF-RIS noise, which is intuitive and also, can be observed from Proposition 1.
Figure 1: (a) Impact of $\phi$ on $\rm SR$}, (b) Impact of noise power at MF-RIS on $\rm SR$}, and (c) Impact of system parameters on the ${\rm SR}-P_{\rm EH}$ region.
@@
Fig. 1(c) illustrates the proposed characterization with the considered system parameters and also, spreading factor $\beta=60$. Specifically, for this set of system parameters, we vary the reference length $\phi$ in $1,\beta]$ such that $\zeta=\frac{\beta}{\phi}\in\mathbb{Z}$ and investigate its combined impact on ${\rm SR}_{t},{\rm SR}_{r},$ and $P_{\rm EH}$, respectively. We observe that, for a given $(N_{t},N_{r},N_{h})$ combination, having $\Upsilon_{t}>1$ and $\Upsilon_{r}>1$ results in significant enhancement of the proposed generalized ${\rm SR}-P_{\rm EH}$ region. Moreover, the figure demonstrates the inevitable role of transmit waveform design on the system performance. Conventional DCSK, i.e., $\zeta=1$ (leftmost points in all the plots) results in the best IT and IR performance but a poor EH performance, which implies a very high $N_{h}^{\min}$ in (IV). On the contrary, $\zeta=\beta$ results in a completely opposite scenario, i.e., worst IT and IR performance but excellent EH performance. Hence, DCSK is not always the best option and depending on the application specific acceptable BER, we decide on $\phi$ for a given $\beta$.
@@
Therefore, in this work, we proposed a DCSK-based MF-RIS-aided noncoherent SWIPT framework. While the IT and IR section of the MF-RIS serve the users in the transmission and reflection region, respectively, the EH section harvests energy from the incident signal to meet their energy requirement. Accordingly, we derived the analytical expression for the system BER and characterized the harvested energy-BER trade-off. We also investigated the impact of this trade-off on the problem of transmit waveform design.
@@
To evaluate ${\rm BER}_{X}$, we need to obtain the mean and variance of $\lambda_{p}^{(X)}$ from (III-A) [7]. Accordingly, by using $\beta=\zeta\phi$ and the definition $E_{b}=P_{\rm t}T_{c}\left(\beta+\phi\right)\mathbb{E}\{x^{2}\}$, we get
@@
$\mathbb{E}\{\lambda_{p}^{(X)}\}$ $=\frac{\beta\gamma_{X,0}}{\left(\beta+\phi\right)}N_{X,0}d_{p}\Lambda_{X,1}\quad{\rm and}$
${\rm var}\{\lambda_{p}^{(X)}\}$ $=\frac{\beta N_{X,0}^{2}}{2}\left(\Lambda_{X,1}\frac{\gamma_{X,0}\left(\zeta+1\right)}{\beta+\phi}+\frac{1}{2}\right)$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!+\frac{\beta\Psi_{X}^{2}N_{0}N_{X,0}}{4}\Lambda_{X,2}\left(\left(\zeta+1\right)+\frac{4\zeta\gamma_{X,0}}{\left(\beta+\phi\right)}\Lambda_{X,1}\right),$ (16)
@@
where $\mathbb{E}\{\cdot\}$ denotes the expectation operatorn. Note that, both $\lambda_{p}|(d_{p}=+1)$ and $\lambda_{p}|(d_{p}=-1)$ are sum of a significantly large number of random variables. Therefore, by using the Gaussian approximation [5] and assuming equally probable transmission of $d_{p}=\pm 1$, the system IT/IR BER based on the channel conditions is obtained as
@@
By replacing (V) in (V) and followed by generalization over the considered channel model, we obtain
${\rm BER}_{X}\!\!=\!\!\!\!\int\limits_{0}^{\infty}\!\!\int\limits_{0}^{\infty}\!\!{\rm BER}_{X}\!\left(\Lambda_{X,1},\Lambda_{X,2}\right)\!f(\Lambda_{X,1})\!f(\Lambda_{X,2})d\Lambda_{X,1}d\Lambda_{X,2}.$